# AI-enhanced Collective Intelligence: The State of the Art and Prospects

### v2

@@ 4.1. Communication
For effective communication in human-AI teamwork, AI must first be able to model human information comprehension accurately and then effectively communicate in a manner understandable to humans. Current AI technologies face limitations in verbal and contextual understanding (Stowers et al., 2021) and cognitive capabilities, such as reasoning about others’ mental states and intentions (Matthews et al., 2021). Detection, interpretation, and reasoning about social cues from a human perspective is imperative to ensure effective coordination but has yet to be achieved by AI. Although, recent developments in LLMs appear promising, further advancing our understanding of how LLMs can be utilized in such a way is crucial (Lu et al., 2023; Schaeffer et al., 2024).
@@
AI’s inability to interpret nonverbal cues and limited self-explanation hinder human collaboration. Addressing these challenges, research suggests the development of AI agents with cognitive architectures that can facilitate both a Machine Theory of the Human Mind (MToHM) and a Human Theory of the Machine Mind (HToMM) will be especially important for supporting the emergence of collective intelligence (Gupta et al., 2023).
4.2. Trust
@@
In team success, a human member’s trust in a machine is pivotal (Nam and Lyons, 2020). Research indicates that users interact differently with AI than humans, often exhibiting more openness and self-disclosure with human partners (Mou and Xu, 2017). Developing explanatory methods that foster appropriate trust in AI is challenging yet crucial for enhancing performance in human-AI teams (Bansal et al., 2021). Meanwhile, over-reliance on AI, especially when faced with incorrect AI advice, can be detrimental and lead to human skill degradation. Hence, it is essential for humans to judiciously rely on AI when appropriate and exercise self-reliance in the face of inaccurate AI guidance (Schemmer et al., 2022), especially in high-stakes situations. More tech literacy, public understanding of AI, and advancements in the technology itself will help in overcoming these challenges.
4.3. Crowd retention
@@
Albeit beneficial in terms of efficiency, employing AI and citizen scientists together has been reported to damage the retention of citizen scientists (Bowyer et al., 2015; Trouille et al., 2019). There is a trade-off between efficiency and retention of the crowds when deploying AI in citizen science projects, and ultimately, if all the crowds leave the projects, the performance of AI alone would decline again. One of the factors contributing to the retention of volunteers is reported to be the social ties they make in their teams (Hristova et al., 2013). The challenge of forming social bonds with AI could increase human feelings of loneliness and isolation, potentially affecting mental health. Questions about identity and recognition could arise when humans are rewarded in conjunction with AI. Studies highlight concerns regarding the anthropomorphic appearance of social robots potentially threatening human distinctiveness (Ferrari et al., 2016). Additionally, collaboration with AI may demotivate humans due to a lack of competitive drive and an over-reliance on AI, which could diminish human participation and initiative. Understanding how to balance AI collaboration with maintaining human motivation and retention is crucial.
4.4. Technology design
@@
The interface design of AI technology influences user experiences, engagement, and acceptance of AI-driven communication. Interfaces should be user-friendly, promoting seamless information exchange and aligning with human cognitive processes. Since AI systems need time to process human input and generate output, there may be a significant time lag between the interactions in some instances. The delay of feedback between human input and AI responses can adversely affect (MacKenzie and Ware, 1993) the coordination and interaction efficiency between humans and AI, especially in collaborative tasks. On the other hand, human behavior possesses specific temporal patterns, mainly determined by circadian rhythm (Yasseri et al., 2012a). It is essential to design such hybrid systems so that the temporal characteristics of both entities match, support mutual understanding, and balance the cognitive load and workload distribution. Moreover, these systems should be designed to augment human contributions (Jarrahi, 2018).
4.5. Ethical considerations
@@
When integrating AI into human interactions, several ethical considerations emerge. Addressing these concerns involves balancing benefits and minimizing potential harms. Key issues include the ethical implications of mandatory AI use, especially when humans are reluctant or do not require AI assistance. When it comes to AI transparency, a study reported that humans who were paired with bots to play a repeated prisoner’s dilemma game outperformed human-human teams initially but later started to defect once they became aware that they were playing with a bot, showing a trade-off between efficiency and transparency (Crandall et al., 2018). Another critical aspect is addressing biases inherent in AI outputs, which are influenced by the data inputs. It is essential to consider how AI integration might exacerbate or introduce new biases in decision-making (Silberg and Manyika, 2019). Furthermore, liability questions arise when AI errors cause serious harm, an issue compounded by the lack of comprehensive regulations and laws. For instance, a robot failed in object recognition and crushed a human worker to death in a factory in South Korea (BBC News, 2023). This highlights the urgent need for comprehensive legal frameworks to address such liabilities and ensure safety in environments where AI interacts with humans. As AI becomes more prevalent in the workplace, constructing ethical human-AI teams becomes a complex, evolving challenge (Flathmann et al., 2021).
4.6. Conclusion
@@
In current real-world AI-CI applications, AI is predominantly utilized as a technical tool to facilitate data analysis through techniques such as natural language processing and computer vision. Implementing AI not only scales up tasks but, in combination with human intelligence, leads to superior performance. Although AI is critical in some decision-making processes, humans make the ultimate decisions and contributions. It is essential to recognize that AI’s function is to support and enhance human collaborative processes rather than to replace human intelligence. Nevertheless, it is also worth considering collective intelligence emerging from a group of AI entities in future research. Past work has demonstrated that the collective behavior of autonomous machines can lead to the emergence of complex social behavior (Tsvetkova et al., 2017a).
@@
AI-CI approaches present numerous challenges. Success in this field relies on multiple factors beyond the initial idea and readiness of technological solutions. Critical aspects include attracting the audience, designing user-friendly interfaces, and effectively integrating AI with human collective intelligence. Understanding how to combine human intelligence and AI to address social challenges remains a significant and worthy area of study.
To better understand the collectives of humans and AI to help overcome societal challenges, several researchers propose to build the foundations of Social AI (Pedreschi et al., 2023) integrating insights from Complex Systems, Network Science, and AI. We agree that researchers across all relevant disciplines should collaborate and keep up with the latest developments to foster a promising future of AI-enhanced collective intelligence.
Acknowledgements.

### v3

@@ Introduction
For effective communication in human-AI teamwork, AI must first be able to model human information comprehension accurately and then effectively communicate in a manner understandable to humans. Current AI technologies face limitations in verbal and contextual understanding 233 and cognitive capabilities, such as reasoning about others’ mental states and intentions 234. Detection, interpretation, and reasoning about social cues from a human perspective is imperative to ensure effective coordination, but AI has yet to achieve this. Although recent developments in LLMs appear promising, further advancing our understanding of how LLMs can be utilized in such a way is crucial 235, 236.
@@
AI’s inability to interpret nonverbal cues and limited self-explanation hinder human collaboration. Addressing these challenges, research suggests the development of AI agents with cognitive architectures that can facilitate both a Machine Theory of the Human Mind (MToHM) and a Human Theory of the Machine Mind (HToMM) will be especially important for supporting the emergence of collective intelligence 168.
Trust
@@
In team success, a human member’s trust in a machine is pivotal 237. Research indicates that users interact differently with AI than humans, often exhibiting more openness and self-disclosure with human partners 238. Developing explanatory methods that foster appropriate trust in AI is challenging yet crucial for enhancing performance in human-AI teams 178. Meanwhile, over-reliance on AI, especially when faced with incorrect AI advice, can be detrimental and lead to human skill degradation. Hence, it is essential for humans to judiciously rely on AI when appropriate and exercise self-reliance in the face of inaccurate AI guidance 239, especially in high-stakes situations. More tech literacy, public understanding of AI, and advancements in the technology itself will help in overcoming these challenges.
Crowd retention
@@
Albeit beneficial in terms of efficiency, employing AI and citizen scientists together has been reported to damage the retention of citizen scientists 240, 206. There is a trade-off between efficiency and retention of the crowds when deploying AI in citizen science projects, and ultimately, if all the crowds leave the projects, the performance of AI alone would decline again. One of the factors contributing to the retention of volunteers is reported to be the social ties they make in their teams 241. The challenge of forming social bonds with AI could increase human feelings of loneliness and isolation, potentially affecting mental health. Questions about identity and recognition could arise when humans are rewarded in conjunction with AI. Studies highlight concerns regarding the anthropomorphic appearance of social robots potentially threatening human distinctiveness 242. Additionally, collaboration with AI may demotivate humans due to a lack of competitive drive and an over-reliance on AI, which could diminish human participation and initiative. Understanding how to balance AI collaboration with maintaining human motivation and retention is crucial.
Technology design
@@
The interface design of AI technology influences user experiences, engagement, and acceptance of AI-driven communication. Interfaces should be user-friendly, promoting seamless information exchange and aligning with human cognitive processes. Since AI systems need time to process human input and generate output, there may be a significant time lag between the interactions in some instances. The delay of feedback between human input and AI responses can adversely affect 243 the coordination and interaction efficiency between humans and AI, especially in collaborative tasks. On the other hand, human behavior possesses specific temporal patterns, mainly determined by circadian rhythm 244. It is essential to design such hybrid systems so that the temporal characteristics of both entities match, support mutual understanding, and balance the cognitive load and workload distribution. Moreover, these systems should be designed to augment human contributions 245. To effectively cater to diverse user groups, human-centered AI design approaches, including the use of personas 246, can be employed to develop adaptive and inclusive human-AI interfaces.
Ethical considerations
@@
When integrating AI into human interactions, several ethical considerations emerge. Addressing these concerns involves balancing benefits and minimizing potential harms. Key issues include the ethical implications of mandatory AI use, especially when humans are reluctant or do not require AI assistance. When it comes to AI transparency, a study reported that humans who were paired with bots to play a repeated prisoner’s dilemma game outperformed human-human teams initially but later started to defect once they became aware that they were playing with a bot, showing a trade-off between efficiency and transparency 247. Another critical aspect is addressing biases inherent in AI outputs, which are influenced by the data inputs. It is essential to consider how AI integration might exacerbate or introduce new biases in decision-making 229. Furthermore, liability questions arise when AI errors cause serious harm, an issue compounded by the lack of comprehensive regulations and laws. For instance, a robot failed in object recognition and crushed a human worker to death in a factory in South Korea 248. This highlights the urgent need for comprehensive legal frameworks to address such liabilities and ensure safety in environments where AI interacts with humans. As AI becomes more prevalent in the workplace, constructing ethical human-AI teams becomes a complex, evolving challenge 249.
Conclusion
@@
In current real-world AI-CI applications, AI is predominantly utilized as a technical tool to facilitate data analysis through techniques such as natural language processing and computer vision. Implementing AI not only scales up tasks but, in combination with human intelligence, leads to superior performance. Although AI is critical in some decision-making processes, humans make the ultimate decisions and contributions. It is essential to recognize that AI’s function is to support and enhance human collaborative processes rather than to replace human intelligence. Nevertheless, it is also worth considering collective intelligence emerging from a group of AI entities in future research. Past work has demonstrated that the collective behavior of autonomous machines can lead to the emergence of complex social behavior 250.
@@
AI-CI approaches present numerous challenges. Success in this field relies on multiple factors beyond the initial idea and readiness of technological solutions. Critical aspects include attracting the audience, designing user-friendly interfaces, scalability, and effectively integrating AI with human collective intelligence. Understanding how to combine human intelligence and AI to address social challenges remains a significant and worthy area of study that requires interdisciplinary collaboration.
While interdisciplinary collaboration presents challenges, such as differences in terminology, methodologies, and objectives across fields, which can create communication barriers and slow progress, it also offers benefits. Interdisciplinary collaboration fosters a comprehensive understanding of complex problems and enables developing more effective, adaptable, and ethical AI-enhanced collective intelligence systems.
Complex system thinking has illuminated various biological, physical, and social domains. Based on this, we can model real-world collective intelligence systems as multilayer networks. This framework allows us to leverage extensive research on multilayer networks, focusing on their robustness, adaptivity, scalability, resilience, and interoperability.
Our theoretical model of human-AI collective intelligence, a multilayer network with cognition, physical, and information layers, offers valuable insights into complex phenomena in such hybrid systems. For example, in the case of the Flash Crash 2012 251, the cognition layer includes high-frequency traders (algorithms) responding to market signals, the physical layer involves the actual financial transactions, and the information layer encompasses the data streams and communication networks connecting these elements. By understanding the interplay and feedback loops within these layers, our model helps identify vulnerabilities, such as how a single large sell order propagated through the network, causing a cascade of automated responses that led to the crash 218.
Similarly, regarding the reported issue of human user retention while AI is being deployed in citizen science projects 206, our model elucidates the interaction between human volunteers and AI classifiers within the different layers. It highlights the importance of maintaining a balance between efficiency (enhanced by AI) and volunteer engagement (rooted in human cognition). This approach prepares us to preemptively address potential pitfalls, ensuring more resilient and effective collective intelligence systems in the future.
Future research directions
Future research in this field will require strong interdisciplinary communication between different domains and research fields. Here are a few directions for future research:
Behavioral Studies
Various behavioral aspects, including human behavior, AI behavior, and the interactions between humans and AI, must be studied. To better understand and harness human-AI collectives for addressing societal challenges, researchers suggest building the foundations of Social AI 48 by integrating insights from Complex Systems, Network Science, and AI.
Adaptive Systems
Adaptive systems can learn and adapt over time. Researchers propose studying collective adaptation 64, focusing on how human collectives adjust their cognitive strategies and social networks in response to changing problems. Extending this research to include both human and AI collective adaptation could be valuable, requiring collaboration from cognitive science, social psychology, and other related fields.
Ethical guidelines
Developing ethical guidelines and frameworks is vital to ensuring the responsible and equitable use of AI-enhanced systems and mitigating potential risks. This includes studying ethics and bias and examining broader societal implications, such as effects on employment, privacy, and social equity.
Future research should focus on large-scale empirical studies to validate theoretical models, understand real-world implications, and refine the deployment of these systems. Interdisciplinary collaborations are crucial for advancing the understanding and development of AI-enhanced collective intelligence systems. We agree that researchers across all relevant disciplines should collaborate and keep up with the latest developments to foster a promising future of AI-enhanced collective intelligence.
Resource availability
Data and code availability
Our source data and code are available on GitHub: https://github.com/cuihaosabrina/AI-enhanced-collective-intelligence.
Supplemental information
Table S1. Main concepts of key terms in human-AI collective intelligence.
Table S2. Summary of CI and AI aspects of various application examples.
Acknowledgments
@@
Author contributions
HC and TY conceived the idea. HC collected the data and performed the analysis. Both authors drafted the manuscript, read and approved the final manuscript.
Declaration of interests
The authors declare no competing interests.